The chlorine present in PVC has two stable isotopes. 35Cl with a mass of 34.97 amu and 37CI with a mass of 36.95 amu. Given chlorine's atomic mass of 35.45, what is the percent abundance of each isotope?
1
Expert's answer
2020-09-30T07:50:07-0400

The atomic mass "A" of chlorine is the weighted average of the atomic masses "A_1 = 34.97" and "A_2=36.95" of its isotope. The isotope masses are multiplied by their abundances "f_1" and "f_2" (in fraction):

"A = f_1A_1 + f_2A_2" .

As there are only two stable isotopes, then, inevitably,

"f_1 = 1-f_2" .

Substituting this into the first equation:

"A = (1-f_2)A_1 + f_2A_2" .

Therefore, the abundance of the second isotope is:

"f_2 = \frac{A-A_1}{A_2-A_1} = \frac{35.45-34.97}{36.95-34.97} = 0.2424" , or 24.24%.

Finally, the abundance of the first isotope is:

"f_1 = 1-f_2 = 1-0.2424 = 0.7576", or 75.76%.

Answer: The percent abundances of the isotopes are 75.76%(34.95 amu) and 24.24%(36.95 amu).